# Spherical Double K-Means: a co-clustering approach for text data analysis

### v2

@@ 4 Cluster validity and simulation study
The simulation study has been developed in order to test the model and algorithm’s performance. We implemented a data generation procedure in order to generate datasets whose clustering structure follows a SDKM model. The generated datasets are then used to focus on different task: (i) development of the pseudo-$F$ index, proposed by Vichi (2015), to detect the true number of clusters of units $K$ and the true number of clusters of variables $Q$; (ii) implementation of different scenarios (corresponding to different error levels) in order to assess model performance in terms of true partitions recovery, measured with the Adjusted Rand Index (ARI) (Hubert & Arabie, 1985), and centroids matrix recovery, measured with Rooted Mean Squared Error (RMSE) and its Normalized versions (NRMSE1, NRMSE2). Clearly, there are several ways to normalize the RMSE: we decided to implement NRMSE1, where the RMSE is normalized by dividing it by the range of the true centroids matrix; NRMSE2, instead, is obtained by dividing RMSE by the norm of centered centroids matrix.
@@ 6.2 Word Clusters: Thematic Comparisons
Beyond their broad similarity in document partitioning, DKM and SDKM exhibit noteworthy differences at the word level. To illustrate how each method groups the main words, we identify the most frequent terms in each cluster, calculate pairwise Jaccard distances among these word sets, and then apply classical multidimensional scaling (MDS).
Label the three DKM clusters as DKM_1, DKM_2, DKM_3 and the three SDKM clusters as SDKM_1, SDKM_2, SDKM_3. For each cluster, we measure the overlap of the main words using the Jaccard distance:
$d_{\text{Jaccard}}(A,B)\;=\;1\;-\;\frac{|A\cap B|}{|A\cup B|},$
where $A$ and $B$ are sets of top words. A smaller value indicates stronger overlap in membership. The resulting $3\times 3$ dissimilarity matrix appears in Table 5.
Table 5: Jaccard dissimilarities among the three clusters of DKM (rows) and SDKM (columns). Lower values imply greater overlap of top words.
SDKM_1 SDKM_2 SDKM_3
DKM_1 0.095 1 0.947
DKM_2 1 0.462 0.824
DKM_3 0.974 1 0.947
Figure 8 then shows the MDS projection of these distances into a 2D space. Clusters that share more words appear closer, while those sharing fewer words appear farther apart. One cluster, centered on words such as “america,” “freedom,” “democracy,” and “dream,”, emerges as quite similar in both methods, whereas the other two clusters diverge more substantially.
Figure 8: MDS plot of word clusters for DKM (green) and SDKM (orange).
Both DKM and SDKM isolate a cluster of aspirational or national-identity terms, including “america,” “freedom,” “democracy,” “child,” “peace,” and “dream.” These words typically co-occur in modern addresses (post Civil War and throughout the 20th century) emphasizing a forward-looking vision for the nation. Because such vocabulary dominates many speeches from that era, Euclidean and cosine-based approaches alike identify a similar “American Dream” cluster.
The main discrepancies arise in how each method groups governance, policy, and constitutional terms. In the SDKM application, for instance, older addresses (1789–1861) focus heavily on “union,” “constitution,” “secession,” and “law,” while later ones shift toward “dream,” “freedom,” “economics,” and “America’s ideal”. Although DKM also detects an older vs. newer thematic line, it often merges older constitutional references with modern economic terms if they appear together in certain large documents. By contrast, SDKM, emphasizing relative usage, splits them if they differ in proportional frequency. For example, words like “race,” “tax,” and “business” are grouped together by DKM if they frequently co-occur in a handful of lengthy speeches, whereas SDKM may place them in separate clusters if they appear in different ratios across those speeches.
These patterns underscore how SDKM’s cosine-based metric isolates specific rhetorical proportions, creating distinct clusters for governance/constitutional issues versus modern policy/economic themes. Meanwhile, DKM lumps some topics together whenever they coexist in the same documents, downplaying subtle proportional distinctions. This aligns with the document-level discrepancy observed in Lincoln’s 1865 speech: a borderline case, rich in “union” and “heal” language, can shift clusters under SDKM, just as certain policy or war-related terms shift between the two methods depending on how they recur in the corpus.
@@ 7 Final considerations
In this paper, we introduced the Spherical Double K-Means (SDKM) clustering method, a novel approach for the simultaneous partitioning of terms and documents in textual data. By integrating the principles of k-means, double k-means, and spherical k-means clustering, SDKM effectively addresses the challenges posed by high dimensionality, sparsity, and noise inherent in text data. The incorporation of cosine similarity within the co-clustering framework allows for a deep understanding of the text and a more meaningful clustering by considering the angular relationships between data points, which is particularly beneficial for textual data where document lengths and term frequencies can vary significantly.
Applying SDKM to the US presidential inaugural addresses dataset demonstrated its capability to uncover distinct thematic clusters that evolved over time. The method not only identified clear chronological distinctions in the clusters of documents but also revealed how historical events have influenced the themes and language used in presidential discourse. For instance, the transition between clusters aligns with significant periods such as the civil war, the great depression and world wars, highlighting shifts in national priorities and rhetoric. This application underscores the potential of SDKM to extract meaningful patterns from complex text corpora, making it a valuable tool for tasks such as topic modeling, sentiment analysis, and information retrieval.
In addition to demonstrating SDKM’s effectiveness on the presidential addresses corpus, we compared its results with Double K-Means using the same TF-IDF weighted data. While both methods identified a broadly consistent chronological split for most documents, one notable discrepancy emerged with Lincoln’s 1865 speech, which SDKM grouped among more modern addresses. This reflects SDKM’s reliance on relative term usage rather than raw vector magnitude, often leading to more thematically cohesive clusters. At the word level, the methods agreed on an “American Dream” cluster but diverged significantly on governance and policy related terms, indicating that SDKM’s cosine‐based co-clustering can isolate finer proportional distinctions. These findings reinforce SDKM’s potential to uncover nuanced patterns in complex textual corpora, complementing more traditional Euclidean-based approaches.

### v3

@@ 4 Cluster validity and simulation study
The simulation study has been developed in order to test the model and algorithm’s performance. We implemented a data generation procedure in order to generate datasets whose clustering structure follows a SDKM model. The generated datasets are then used to focus on different task: (i) development of the pseudo-$F$ index, proposed by Vichi (2015), to detect the true number of clusters of units $K$ and the true number of clusters of variables $Q$; (ii) implementation of different scenarios (corresponding to different error levels) in order to assess model performance in terms of true partitions recovery, measured with the Adjusted Rand Index (ARI) (Hubert & Arabie, 1985), and centroids matrix recovery, measured with Rooted Mean Squared Error (RMSE) and its Normalized versions (NRMSE1, NRMSE2). Clearly, there exist several normalization techniques: we decided to implement NRMSE1, where the RMSE is normalized by dividing it by the range of the true centroids matrix; NRMSE2, instead, is obtained by dividing RMSE by the norm of centered centroids matrix.
@@ 6.2 Word Clusters: Thematic Comparisons
Beyond their broad similarity in document partitioning, DKM and SDKM exhibit noteworthy differences at the word level. To illustrate how each method groups the main words, we build a contingency table, represented in Table 5, that compares the cluster assignments for all words. Specifically, each row corresponds to a DKM cluster, and each column to a SDKM cluster.
Table 5: Contingency table for all words assigned to each pair of clusters (DKM rows, SDKM columns).
[…]
1 201 0 0
2 0 295 143
3 256 125 185
The Table reveals how words are distributed across the six possible cluster pairs. For example, 201 words appear in both cluster 1 under DKM and cluster 1 under SDKM, whereas 256 words assigned to DKM 3 fall into SDKM 1. These differences reflect how each method partitions the same vocabulary, highlighting areas of overlap and divergence. To further investigate differences in the two partitions, we use the ARI to measure the similarity between two partitions; it usually ranges in $[0,1]$ and it is equal to 1 when the two partitions perfectly match, while it is equal to 0 when one of the two partitions can be considered as a partition obtained by chance. The two partitions are dissimilar, as also the ARI index equal to 0.225 reveals.
By exploring differences in the clustering interpretation, we considered the ”top words” (words with the 30 highest frequencies) in each cluster. In fact, these words help to label each cluster with a specific topic.
Figure 8 provides a complementary view of these differences. In each panel, we compare the top words identified by DKM and by SDKM for the same cluster index. If a word appears in both methods’ top 30, it is shown in purple (“Overlap”), while words unique to DKM or SDKM are colored differently.
Figure 8: Top words in each cluster for DKM vs. SDKM. Purple bars indicate overlap in both methods, green indicates DKM only, and orange indicates SDKM only.
Both DKM and SDKM isolate a cluster of aspirational or national-identity terms, including “america,” “freedom,” “democracy,” “child,” “peace,” and “dream.” These words typically co-occur in modern addresses (post–Civil War and throughout the 20th century), emphasizing a forward-looking vision for the nation. Because such vocabulary dominates many speeches from that era, both Euclidean- and cosine-based approaches identify a similar “American Dream” cluster, which is apparent in Figure 8 through the substantial overlap (purple bars) in the first panel. The second cluster, highlighted by words like “union,” “constitution,” “state,” and “revenue,” centers on governance and legal structures. These foundational words appear mostly as overlap, reflecting how both methods capture a “Law and Order” theme. Nevertheless, DKM alone includes terms such as “congress” and “law” in its top list, whereas SDKM alone brings in additional administrative or procedural vocabulary such as “case” and “department.” This suggests that SDKM’s emphasis on relative usage encourages a slightly finer partitioning of specialized policy references, whereas DKM merges them when they appear together in large documents. The third cluster, focusing on politics, economic considerations, and older secession-related rhetoric, shows somewhat greater divergence. Figure 8 reveals that SDKM uniquely elevates words like “negro,” “tariff,” and “amendment,” while DKM includes “community,” “control,” and “progress” in the same top set. Although certain vocabulary, such as “race” and “congress,” again appears in both methods’ top words, the larger number of unique terms indicates that SDKM splits off certain policy and historical references more decisively, presumably by tracking proportional usage across speeches. DKM, on the other hand, merges them whenever they tend to co-occur in lengthier addresses, thereby placing broad economic and societal concepts together. These patterns underscore how SDKM’s cosine-based metric isolates specific rhetorical proportions, creating distinct clusters for governance/constitutional issues versus modern policy/economic themes. Meanwhile, DKM groups some topics together whenever they coexist in the same documents, downplaying subtle proportional distinctions. This aligns with the document-level discrepancy observed in Lincoln’s 1865 speech: a borderline case, rich in “union” and “heal” language, can shift clusters under SDKM, just as certain policy or war-related terms shift between the two methods depending on how they recur in the corpus.
@@ 7 Final considerations
In this paper, we introduced the Spherical Double K-Means clustering method, a novel approach for the simultaneous partitioning of terms and documents in textual data. By integrating the principles of K-Means, Double K-Means, and Spherical K-Means clustering, SDKM effectively addresses the challenges posed by high dimensionality, sparsity, and noise inherent in text data. The incorporation of cosine similarity within the co-clustering framework allows for a deep understanding of the text and a more meaningful clustering by considering the angular relationships between data points, which is particularly beneficial for textual data where document lengths and term frequencies can vary significantly. In fact, traditional k-means using Euclidean distance often produces large, dominant clusters alongside smaller residual ones when applied to textual data, even after TF-IDF weighting. This tendency occurs because Euclidean distance is most effective when variables are uncorrelated, a condition seldom met in language data where terms frequently co-occur in highly correlated patterns. By contrast, SDKM, employing cosine similarity, focuses on the relative distribution of term usage rather than absolute frequencies. In doing so, it captures the angular relationships between document vectors, thereby disentangling topics more effectively and yielding clusters that are more semantically coherent. Applying SDKM to the US presidential inaugural addresses dataset demonstrated its capability to uncover distinct thematic clusters that evolved over time. The method not only identified clear chronological distinctions in the clusters of documents but also revealed how historical events have influenced the themes and language used in presidential discourse. For instance, the transition between clusters aligns with significant periods such as the civil war, the great depression and world wars, highlighting shifts in national priorities and rhetoric. This application underscores the potential of SDKM to extract meaningful patterns from complex text corpora, making it a valuable tool for tasks such as topic modeling, sentiment analysis, and information retrieval.
In addition to demonstrating SDKM’s effectiveness on the presidential addresses corpus, we compared its results with Double K-Means using the same TF-IDF weighted data. While both methods identified a broadly consistent chronological split for most documents, one notable discrepancy emerged with Lincoln’s 1865 speech, which SDKM grouped among more modern addresses. At the word level, the methods agreed on an “American Dream” cluster but diverged significantly on governance and policy related terms, indicating that SDKM’s cosine‐based co-clustering can isolate finer proportional distinctions. These findings reinforce SDKM’s potential to uncover nuanced patterns in complex textual corpora, complementing more traditional Euclidean-based approaches.